# Target Aggregate Data Adjustment Method for Transportability Analysis Utilizing Summary-Level Data from the Target Population

### v1

@@ 5 Results
Table 5: The simulation results of scenarios 1 to 10 when the overall censoring proportion is around 50%. The columns from left to right represent scenario index; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo true hazard ratio; coverage of pseudo true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo true hazard ratio when adopting censoring-ignored strategy.
@@ 6 Discussion
In this paper, we demonstrated that the proposed TADA method provides substantial improvements in bias control and coverage across a range of censoring scenarios, showing particular effectiveness under low to moderate censoring proportions. These findings support the value of TADA in addressing censoring bias, with its advantage over unadjusted methods persisting even as overall censoring increases, albeit with some reduction in performance. Overall, these results suggest that TADA structurally addresses information loss due to censoring, making it a robust transportability analysis method in scenarios where IPDs are unavailable for the target population and where source data are disturbed by right-censoring.
@@
A core strength of the proposed TADA method lies in its flexible weighting by a product structure, with each term specifically designed to adjust for different sources of confounding. This makes TADA highly adaptable to complex application scenarios. In current cases in TADA, the participation weights, derived via the MoM, can address the challenges of working with AgD for the target population for transportability analysis. In this study, we explored TADA’s ability to transport survival outcomes under covariant-relevant censoring, which addresses the current gap in the transportability methods that generally have been limited to binary and continuous outcomes with requirements of IPD.
@@
In scenarios with modest censoring rates, the TADA method with an adjustment for censoring exhibits robust performance with low bias and high coverage. This effectiveness can be attributed to the lower level of information loss in low-censoring contexts, allowing TADA’s inverse probability weighting to make more precise adjustments. The stability of TADA’s performance in these scenarios highlights its ability to achieve accurate effect estimates under conditions of limited censoring, where the method leverages available data efficiently. In scenarios with a higher overall censoring proportion, TADA demonstrates performance close to its practical limit, with a noticeable decline in bias control and coverage. Increased censoring results in a higher loss of information that in turn effectively reduces the sample size available for reliable adjustments. Higher censoring may result in more extreme weights and introduce greater variability in bias control. This accumulation of censoring-induced bias and weight instability creates challenges for maintaining accuracy and coverage under high-censoring conditions. Nonetheless, censoring-adjusted TADA outperforms the unadjusted approach in general.
@@
In scenarios where there was a larger disparity in censoring proportions between treatment and control groups, TADA shows lower bias estimates. This phenomenon suggests that TADA may potentially benefit from this unique data characteristic. When one group has lower censoring, the more complete data could allow TADA’s weighting mechanism to draw on relatively stable information, which may enhance the reliability of adjustments. This natural heterogeneity might provide a form of implicit stratification, helping TADA balance distinct group effects with greater robustness. Such conditions could, in theory, offer additional stability in TADA’s performance by leveraging the more consistent information from less-censored data. However, as overall censoring increases, this differential effect diminishes, indicating that higher censoring rates may challenge TADA’s capacity to control bias across varying group censoring proportions.
Future research could expand the scope of TADA’s applicability by exploring alternative approaches to deriving weights for population balance and handling censoring adjustments. In the current framework, the MoM is employed to achieve population balance, while IPCW is used to adjust for censoring. However, the effectiveness of IPCW is somewhat reliant on the PH assumption, which may limit its utility in settings where this assumption does not hold. Exploring alternative censoring adjustment techniques that do not depend on the PH assumption could enhance TADA’s versatility in handling varied censoring structures 18. By experimenting with different weighting schemes for censoring adjustments, TADA could potentially expand its applicability in contexts where data characteristics or censoring patterns diverge from the ideal settings. This evolution would allow TADA to maintain its effectiveness under diverse data conditions, making it a more adaptable tool across a broad range of practical applications. Additionally, while TADA is currently applied for RCT source data, extending its framework to accommodate observational data could further broaden its utility, enabling transportability analyses in settings where randomized data is unavailable.
@@ 7 Conclusion
Our study demonstrates the TADA’s robust ability to effectively address censoring bias across a broad spectrum of scenarios, demonstrating superior bias control and coverage even under challenging conditions. The TADA method’s design for using aggregate data for causal inference, along with its ability to accommodate varying censoring proportions, makes it an invaluable tool for transporting survival estimates, especially in settings where target IPD is hard to access.
@@ Acknowledgments
Conflict of Interest: None declared.

### v2

@@ 6.3 Abnormal Values in the Target Population
Table 6: The simulation results of various source sample size $N_{\textit{s}}$ with corresponding Monte Carlo SD and Bootstrap SE. The columns from left to right represent source sample size $N_{\textit{s}}$; bias and 95% CI between censoring-adjusted strategy and pseudo-true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo-true hazard ratio; coverage of pseudo-true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo-true hazard ratio when adopting censoring-ignored strategy; Monte Carlo SD for the censoring-adjusted strategy; Bootstrap SE for the censoring-adjusted strategy; Monte Carlo SD for the censoring-ignored strategy; Bootstrap SE for the censoring-ignored strategy. All the other settings in the simulations are identical. SD Standard deviation. SE Standard error.
ABV Type HR (95% CI): Censor HR (95% CI): Non-censor HR (95% CI): True Bias (95% CI): Censor Bias (95% CI): Non-censor Coverage: Censor Coverage: Non-Censor
1 0.858 (0.575, 1.296) 0.932 (0.625, 1.412) 0.818 (0.811, 0.826) 0.039 (–0.253, 0.437) 0.114 (–0.210, 0.543) 92.80% 91.40%
2 0.858 (0.574, 1.295) 0.932 (0.624, 1.410) 0.818 (0.811, 0.826) 0.039 (–0.251, 0.439) 0.114 (–0.207, 0.559) 93.20% 91.80%
3 0.857 (0.573, 1.294) 0.931 (0.623, 1.408) 0.818 (0.811, 0.826) 0.038 (–0.258, 0.428) 0.113 (–0.206, 0.543) 94.20% 92.00%
4 0.862 (0.561, 1.338) 0.939 (0.611, 1.465) 0.818 (0.811, 0.826) 0.044 (–0.257, 0.452) 0.121 (–0.206, 0.595) 93.80% 92.40%
5 0.881 (0.493, 1.673) 0.971 (0.539, 1.737) 0.818 (0.811, 0.826) 0.063 (–0.320, 0.629) 0.152 (–0.279, 0.791) 95.00% 93.40%
6 0.860 (0.561, 1.335) 0.937 (0.611, 1.463) 0.818 (0.811, 0.826) 0.042 (–0.272, 0.448) 0.119 (–0.229, 0.570) 93.60% 93.20%
7 0.882 (0.493, 1.680) 0.972 (0.535, 1.748) 0.818 (0.811, 0.826) 0.063 (–0.323, 0.664) 0.153 (–0.269, 0.846) 94.80% 93.20%
Table 7: The simulation results of various abnormal value patterns in the target population. The columns from left to right represent abnormal value types in Table 4; estimated hazard ratio and 95% CI when TADA adjusts censoring; estimated hazard ratio and 95% CI when TADA ignores censoring; pseudo-true hazard ratio and 95% CI; bias and 95% CI between censoring-adjusted strategy and pseudo-true hazard ratio; bias and 95% CI between censoring-ignored strategy and pseudo-true hazard ratio; coverage of pseudo-true hazard ratio when adopting censoring-adjusted strategy; coverage of pseudo-true hazard ratio when adopting censoring-ignored strategy. All the other settings in the simulations are identical. ABV Abnormal value.
7 Real Case Study
To demonstrate the real-world applicability and practical value of the proposed TADA method, we conduct an application that transports survival outcomes from an RCT with censoring to a real-world population with only aggregate-level data. The source dataset is derived from the SQUIRE trial (NCT00981058) 23, a phase III, multi-center, open-label RCT evaluating the addition of necitumumab to gemcitabine and cisplatin chemotherapy in patients with stage IV squamous non-small-cell lung cancer (NSCLC). Patients are randomized in a 1:1 ratio to receive either necitumumab plus chemotherapy or chemotherapy alone. We use the publicly available IPD from the Project Data Sphere platform 43, which contained only patients randomized to the control (gemcitabine/cisplatin) arm ($N_{\textit{s}}$ = 548). Available source sample data includes overall survival time, censoring status, and relevant baseline covariates such as age, sex, ethnic origin, smoking history and Eastern Cooperative Oncology Group (ECOG) performance status. The overall censoring proportion is around 19%.
[…]
Table 8: Baseline characteristics of the source population in SQUIRE trial and the target population in Ontario study.
Source (SQUIRE) Target (Ontario)
Sample size 548 2,056
Age (mean in years) 61.7 72.0
Male (%) 83.6 65.1
The estimand of interest is the marginal survival function over time, with particular focus on the estimated median OS and corresponding 95% confidence intervals. The survival function is estimated using weighted Kaplan–Meier methods under various adjustments on covariate imbalancing and censoring. We apply age and sex as the population matching covariates while adding ethnic origin, smoking history and ECOG performance status when estimating the censoring weights. In particular, we consider four adjustment strategies as follows:
  • Original: The standard Kaplan–Meier estimator applied directly to the source IPD, without any weighting, serving as a naive benchmark that ignores both covariate imbalances and censoring-related bias.
  • Original with Censoring Adjusted: An estimator that applies censoring weights to the source data in order to mitigate bias due to incomplete follow-up, but without accounting for population-level covariate mismatches.
  • Transported without Censoring Adjusted: A weighted Kaplan–Meier estimator incorporating participation weights to align covariate distributions between the source and target populations, while disregarding potential bias introduced by censoring.
  • TADA (Transported with Censoring Adjusted): The proposed TADA method combining both participation weights and censoring weights, simultaneously correcting for censoring and aligning the source population with the target population on the key covariates.
To obtain the truncated final weights, we truncate the raw final weights at a preferred threshold quantile based on their distribution after assessments. Figure 3 illustrates this process: for the shown raw final weights, applying truncation at the 95% quantile could be the optimal choice for current scenario to keep most of information and avoid the extreme weight candidates. We apply the 95% quantile as the truncation threshold for the introduced adjustments in the real case study, supported by assessment results.
Figure 3: An illustration of the final weights truncation threshold decision. The horizontal axis represents the value of raw final weights. The vertical axis indicates the percentage of observations in each weight bin, relative to the total sample. Blue bars depict the distribution of raw final weights. Red vertical dashed lines from left to right indicate the candidate thresholds at the $90^{\text{th}}$, $95^{\text{th}}$ and $99^{\text{th}}$ percentiles, respectively.
Figure 4 displays the comparison between the unadjusted naive Kaplan–Meier curve and two various weighted Kaplan–Meier curves, representing the transported cases with and without censoring adjustment respectively. Table 9 reports the corresponding estimated median OS and 95% confidence intervals. Table 10 summarizes the estimated survival probabilities at every 6 months under different adjustments. The original median OS without any adjustment is 9.856 months, while the transported median OS without censoring adjustment is 9.528 months. Adjusting only for censoring leads to a close median OS as the naive case of 9.889 months. The fully adjusted TADA estimator, simultaneously correcting for both covariate imbalance and censoring bias, yields an estimated median OS of 10.579 months and a narrower confidence interval compared to that of the transported without censoring adjustment result. These estimates illustrate the differences arising from adjustments for censoring and covariate distribution imbalance when transporting clinical trial outcomes to real-world populations, which may lead to different clinical conclusions. Seung et al. 24 also emphasized that real-world cohorts often include older patients compared to those enrolled in clinical trials. As such, the ability of TADA to account for population differences and censoring enhances the external validity of the transported estimates. Due to limited data availability, this case study incorporated only the control arm datasets and limited set of covariates. The inclusion of additional information such as the treatment arm data and more key covariates would yield more clinically relevant survival estimates.
Figure 4: Original Kaplan–Meier curve and Transported Kaplan–Meier curves under various adjustments. Top panel: Original naive K-M curve vs. Transported K-M curve that adjusted for covariate imbalance and without censoring adjustment. Bottom panel: Original naive K-M curve vs. Transported K-M curve fully adjusted by TADA, incorporating both censoring and covariate imbalance adjustments. Dashed vertical lines in grey indicate the estimated median OS for the naive curve. Dashed vertical lines in red indicate the estimated median OS for the transported curves.
Table 9: Estimated median overall survival (OS) time and 95% confidence intervals under different adjustments.
Median OS (months) 95% CI (Lower) 95% CI (Upper)
Original 9.856 8.903 11.138
Original (with Censoring Adjusted) 9.889 8.903 11.138
Transported (w/o Censoring Adjusted) 9.528 7.425 11.696
TADA fully adjusted 10.579 8.608 12.485
Table 10: Estimated survival probabilities at every 6 months under different adjustments.
Month Original Original (with Censoring Adjusted) Transported (w/o Censoring Adjusted) TADA
6 0.723 0.723 0.721 0.738
12 0.428 0.428 0.401 0.443
18 0.243 0.242 0.222 0.254
24 0.165 0.165 0.157 0.183
30 0.106 0.106 0.129 0.141
36 0.055 0.056 0.114 0.107
Table 9 and 10 show that, compared with the naive case, the trend of weighted survival curve remains similar when only adjusted with censoring weights. This consistency results from the fact that few observations are censored in the first 9 months, leading to the estimated censoring weights concentrate near 1 and the Kaplan–Meier estimate largely unaffected. Only in the far right tail of the distribution (i.g., beyond 18 months) does censoring become more frequent, at which point censoring weights exert a modest influence on the estimated survival curve. Figure 5 visually demonstrates this phenomenon.
Figure 5: Distribution of overall survival times by event status when only conduct censoring adjustment for the source SQUIRE population. The grey bars represent the status of event. The navy blue bars represents the status of censor. The red dashed line indicates the value of median OS.
@@ 8 Discussion
In this paper, we demonstrate that the proposed TADA method provides improvements in bias control and coverage across a range of censoring scenarios, showing effectiveness under low to moderate censoring proportions. These findings support the value of TADA in addressing censoring bias, with its advantage over unadjusted methods persisting even as overall censoring increases, albeit with some reduction in performance. The case study between SQUIRE and Ontario population provides an illustration of TADA’s value in real application. Overall, these results suggest that TADA structurally contributes handling with information loss due to censoring, making it a valuable transportability analysis method in scenarios where IPDs are unavailable for the target population and where source data are disturbed by right-censoring.
@@
A core advantage of the proposed TADA method lies in its flexible weighting-by-a-product structure, with each term specifically designed to adjust for different sources of confounding. This makes TADA highly adaptable to complex application scenarios. In current cases in TADA, the participation weights, derived via the MoM, can address the challenges of working with AgD for the target population for transportability analysis. In this study, we explored TADA’s ability to transport survival outcomes under covariate-relevant censoring, which addresses the current gap in transportability methods which have generally been limited to binary and continuous outcomes and require IPD.
@@
In scenarios with modest censoring rates, the TADA method with an adjustment for censoring exhibits robust performance with low bias and high coverage. This effectiveness can be attributed to the lower level of information loss in low-censoring contexts, allowing TADA’s inverse probability weighting to make more precise adjustments. The stability of TADA’s performance in these scenarios highlights its ability to achieve accurate effect estimates under conditions of limited censoring, where the method leverages available data efficiently. In scenarios with a higher overall censoring proportion, TADA demonstrates performance close to its practical limit, with a decline in bias control and coverage. Increased censoring results in a higher loss of information which in turn reduces the sample size available for reliable adjustments. Higher censoring may result in more extreme weights and introduce greater variability in bias control. This accumulation of censoring-induced bias and weight instability creates challenges for maintaining accuracy and coverage under high-censoring conditions. Nonetheless, censoring-adjusted TADA outperforms the unadjusted approach in general.
@@
In scenarios where there was a larger disparity in censoring proportions between treatment and control groups, TADA shows lower bias estimates. This phenomenon suggests that TADA may potentially benefit from this unique data characteristic. When one group has lower censoring, the more complete data could allow TADA’s weighting mechanism to draw on relatively stable information, which may enhance the reliability of adjustments. This natural heterogeneity might provide a form of implicit stratification, helping TADA balance distinct group effects with greater robustness. Such conditions could, in theory, offer additional stability in TADA’s performance by leveraging the more consistent information from less-censored data. However, as overall censoring increases, this differential effect diminishes, indicating that higher censoring rates may challenge TADA’s capacity to control bias across varying group censoring proportions. In our simulation study, we adopted the nonparametric bootstrap to estimate standard errors and coverage. This choice was driven by the method’s flexibility and its compatibility with the complex, non-closed-form weighting scheme in TADA. While bootstrapping is widely used in survival analysis and performs well under low to moderate censoring, it may become less reliable when substantial information is lost due to heavy censoring, which could partially explain the reduced coverage observed in some scenarios with 30% or higher censoring. We note, however, that the primary purpose of our simulation study was to evaluate performance trends in bias control and coverage across scenarios, rather than to identify optimal variance estimation procedures under all conditions. Alternative inference strategies, such as robust sandwich variance estimators 44, may offer improved performance under heavy censoring, and represent a potential extension direction for future research.
It is worth mentioning that the TADA framework itself does not require assumptions such as proportional hazards to be valid in the target population. Rather, TADA’s objective is to reweight the study data so that the distribution of covariates aligns with that of the target population, thus facilitating an unbiased estimation of treatment effects at the target-population level. In the current implementation, we rely on the Cox PH model under the PH assumption, allowing us to interpret the TADA estimate as a constant hazard ratio. If non-proportional hazards are present, this Cox-based result may still be viewed as a time-averaged hazard ratio; yet, investigators seeking to capture time-varying effects more accurately could employ alternative outcome models, such as accelerated failure time models 42, 45, piecewise Cox models 46, or additive hazards frameworks 47. In such scenarios, the reweighting scheme of TADA remains applicable. Users simply need to replace the Cox PH with a more suitable outcome model, and TADA will yield a corresponding estimator that better reflects the underlying hazard structure in the target population.
Future research could expand TADA’s scope by investigating additional weighting strategies for achieving population balance and by exploring more flexible techniques for censoring adjustments. Although our current implementation uses the MoM to align study and target populations, and a standard form of IPCW to address censoring, there remains room to adopt censoring models that do not rely on proportional hazards or similarly restrictive assumptions 18. Examining these alternatives could enhance TADA’s versatility in settings where the censoring patterns deviate from idealized conditions. Exploring approaches such as doubly robust or semiparametric methods would also help alleviate partial misspecification of either weighting component. Finally, while TADA has thus far been demonstrated using RCT data, extending it to accommodate other experimental and observational studies offers another avenue for broadening its applicability, particularly in cases where randomized data may be unavailable.
@@ 9 Conclusion
Our study presents TADA’s performance to address censoring bias across various scenarios, demonstrating effective bias control and coverage even under challenging conditions. The TADA method’s design for using aggregate data for causal inference, along with its ability to accommodate varying censoring proportions, makes it a valuable tool for transporting survival estimates, especially in settings where target IPD is may not be available.
@@ Acknowledgments
The authors wish to thank Dr. Daniel Daly Grafstein for his contributions during the revision of the manuscript. The authors also thank the Editor and anonymous referees of this article for their valuable and thoughtful input. These suggestions have significantly enhanced the quality of the manuscript.
Conflict of Interest
The authors declare no conflicts of interest.
Data Availability Statement
All data used in the simulation study were generated according to the data-generating mechanisms specified in the manuscript. The individual-patient level data for the real case study are openly available in the Project Data Sphere dataset at https://doi.org/10.34949/05tc-jx78, reference number NCT00981058.